# Catch Me if You Can: Detecting Unauthorized Data Use in Deep Learning Models

### v1

@@ Abstract
This work proposes MembershipTracker, a practical data provenance tool that can empower ordinary users to take agency in detecting the unauthorized use of their data in training DL models. We view tracing data provenance through the lens of membership inference (MI). MembershipTracker consists of a lightweight data marking component to mark the target data with small and targeted changes, which can be strongly memorized by the model trained on them; and a specialized MI-based verification process to audit whether the model exhibits strong memorization on the target samples.
@@
Overall, MembershipTracker only requires the users to mark a small fraction of data (0.005$\sim$0.1in proportion to the training set), and it enables the users to reliably detect the unauthorized use of their data (average 0% FPR@100% TPR). We show that MembershipTracker is highly effective across various settings, including industry-scale training on the full-size ImageNet-1k dataset. We finally evaluate MembershipTracker under multiple classes of countermeasures.
1 Introduction
@@
Modern deep learning (DL) models have attained remarkable performance in various tasks such as image classification and language generation. Their success is largely driven by the availability of massive training data [52, 50]. However, it is not always clear whether the data used to train these models has been used with the permission of the data owners. For instance, Clearview.ai, in developing their facial recognition system, scraped billions of user photos from social media platform without the users’ consent [26]. The unregulated use of personal data for building DL models has engendered mounting concerns [20, 66, 26], and is in violation of privacy laws such as the European Union’s General Data Protection Regulation (GDPR) [2]. In spite of this, data holders often have limited agency in detecting the unauthorized use of their data due to the lack of practical data provenance tools.
@@
This leads to two main lines of work for data provenance. The first line of work seek to inject designated features into the user data, which can induce the model to exhibit a certain detectable behavior for provenance purposes [53, 27, 39, 68]. The injected features can take the form of a backdoor trigger pattern [27, 39]; or a spurious feature pattern that can cause output probability shift by the model [68]. However, these techniques either impose strong assumptions on the users (e.g., the ability to mark a large fraction of data) [53, 39, 27] and/or suffer from limited auditing performance [53, 39, 68].
@@
The second line of work is based on membership inference (MI). There are instance- and user-level MI - the former infers whether a specific sample is used to train a model [56, 8]; while the latter detects the usage of any of the user’s data [58, 12, 32]. Our work focuses on auditing whether specific user samples are used for training the target model. Unfortunately, existing MI methods often suffer from limited effectiveness, as measured by their true positive rate (TPR) and false positive rate (FPR) [8]. However, a high-power MI method is needed to facilitate data provenance with low FPR (to avoid false accusation) and high TPR (for correct data provenance).
@@
To improve, recent studies propose several techniques based on data poisoning [65, 13, 10]. Their common idea is to inject mislabeled samples into the model’s training set, which can amplify the model’s behavioral changes on some target samples, thereby making them easier to be de-identified. To apply these techniques for data provenance, the users have to be able to mislabel some training data, which, unfortunately, excludes the real-world scenarios where the data labels are assigned by the model owner or an external service [21, 19]. This renders these techniques challenging for real-world use.
@@
Figure 1: MembershipTracker is a data provenance tool that operates by: (1) marking the target data (e.g., facial images, artworks) with small and targeted changes, and then (2) initiating a specialized membership inference process to audit whether the target data are used for training the model.
@@
MembershipTracker operates under the realistic setting where we assume the users have neither the capability nor expertise to train shadow/proxy models [70, 6, 8], or to manipulate the data labels [65], as these are well beyond the capacity of ordinary personal users (Section 3.1.1). Rather, MembershipTracker only requires the users to mark a small fraction of data (amounting to 0.005$\sim$0.1of the training set) for provenance purposes. Our goal is to empower the users to carry out high-power MI on their target data with high TPR and low FPR, given black-box access to the model.
Technical design. We first observe that the effectiveness of MI on a given model is directly related to the model’s ability in memorizing individual data points, and hence samples that are strongly memorized by the model (e.g., outlier samples) are easier to be de-identified [71, 8, 65, 14]. Therefore, in order to fulfill our goal, there are two challenges as follows.
  1. 1.
@@
For the first challenge, we propose a lightweight data marking technique to mark the target data with small and targeted changes, which are formulated as the combination of outlier feature and procedural noise [17]. These subtle changes can be readily created by the users without assuming access to a proxy model, and can still induce the model to strongly memorize the specially-marked samples while preserving their visual information (e.g., Fig. 1).
For the second challenge, given a target user’s small set of marked samples, we propose a novel set-based MI process to audit whether they are used to train the model. Our approach follows a common idea in existing user-level MI studies [58, 47, 32] to leverage the collective information across the small set of target samples for MI. But unlike prior methods, MembershipTracker does not require any additional shadow/reference models [58, 47, 32] and can enable high-power MI for reliable data provenance.
@@
We find that by merely marking a small fraction of samples (0.005$\sim$0.1of the dataset), MembershipTracker effectively empowers the users to trace the provenance of their data with minimal false positives (average 0% FPR@100% TPR), and it is also scalable to the large-scale ImageNet training (on both Convolutional and Transformer network). This renders MembershipTracker a practical data provenance tool and contributes to responsible AI practice.
2 Related Work
Our work focuses on detecting the unauthorized use of personal data in training DL models, which is complementary to existing work that aim to protect the data from unwanted use by rendering them un-learnable [55, 30]. Hence we focus on existing literature for data provenance, and we classify them into two categories.
@@ 2 Related Work
Dataset level provenance aims to detect whether a DL model is trained on a specific dataset [46, 53, 39]. Maini et al. propose dataset inference, which detects whether two models trained on the same dataset exhibit similarities in their decision boundaries  [46]. Other work modify portions of training set to leave a certain artifact on the models, such as radioactive data [53] and backdoor watermark [39]. However, recent work [31] finds that these techniques [53, 39] still have poor auditing performance.
Moreover, dataset level solutions require several assumptions that are outside the scope of our work. First, many of them require control over a nontrivial portion of training set (e.g., 10$\sim$20[53, 39]). Some solutions assume access to some known feature extractor [53], access to the training set or the ability to train a proxy model [46, 53]. These are well beyond the user capability we consider (in Section 3.1.1).
User level provenance seeks to detect whether the users’ data are used to train a DL model [27, 68], and they require only modifying the users’ own data. Their idea is to mark the users’ data with a designated feature to induce a detectable behavior into the model. Hu et al. propose to inject backdoor trigger feature into the user data [27], but they assume the user capability to manipulate the data labels. Other studies on clean-label backdoor attacks may be repurposed to overcome this, but they still require training a proxy model to compute the poisoned data [59, 73].
Wenger et al. propose to inject a target spurious feature into the user data to be learned by the suspected model, and then detect whether the model has associated the injected feature with the target class label [68]. To detect, the users can separately overlay the target and non-target features to some auxiliary data (outside the target class), and then determine whether the target feature causes a slightly higher prediction probability on the target class, compared with the non-target features (e.g., 10% vs. 1%). Both their work and ours involve adding spurious features to mark users’ data, but there are several major differences between the two. Notably, our analysis reveals that they [68] significantly underestimate their detection false positives (from 0% to >30%). This is because they employ a discrepant detection procedure for the testing features that are used for training (true positives), and those that are not (false positives); however, a fair evaluation should follow a consistent procedure (details in Appendix D). Moreover, we also realign their technique for MI-based data provenance in our work, and find that even when their approach has incurred higher distortion to the data, it still has poorer MI performance than MembershipTracker (Appendix D).
@@
Instance-level MI detects whether a specific instance is used to train a model. It was first proposed by Shokri et al. [56] and there are many follow-up studies [71, 16, 6, 70, 8]. However, even state-of-the-art MI methods [8, 6, 70] still struggle with achieving high TPR when controlled at the low FPR regime (more in Section 4.1).
@@
This has led to several attempts at improving the effectiveness of existing MI methods via data poisoning. Tramer et al. propose to inject mislabeled samples into the model’s training set, which can transform the target samples into outliers and amplify their influence on the model’s decision, thereby making the target samples easier to be de-identified [65]. While injecting mislabeled samples represents an effective solution for improving the MI success [65, 13, 10], it may be challenging for the users to apply to their data, because in many scenarios the users may not have control over the data labels [21, 68, 19]. Even if this is feasible, we find that such a method can still only increase the MI success to a limited extent (validated in Appendix E). By contrast, MembershipTracker does not assume control of the data labels, and is still able to facilitate high-power MI for tracing data provenance.
@@
User-level MI aims to detect whether any of a user’s data is used to train a model. Unlike instance-level MI, the auditor possesses a set of samples from a target user that are not necessarily used to train the target model. Song et al. [58] propose the first user-level MI for natural language models, which first trains multiple shadow models, and then uses the outputs from the shadow models as the features to train an audit model. Subsequent work develop solutions under various settings, including speech recognition models [48, 11], metric learning models [38, 12] and large language models [32, 47].
@@
A common thread of existing user-level MI methods is to aggregate the information across multiple samples (e.g., different voices of the same speaker) to perform MI [58, 11, 12, 32]. The MI process in MembershipTracker follows a similar philosophy. However, prior work often requires access to additional reference models [32] or shadow models [58, 12, 11], and they still have limited MI effectiveness (e.g., limited TPR under the low FPR regime [32, 47]). MembershipTracker overcomes these limitations by means of a novel data-marking and specialized MI auditing process, which can simultaneously achieve high TPR and low FPR for data provenance.
@@
Concurrent work by Huang et al. [31] proposes a general framework for auditing unauthorized data use in training DL models. Their idea is to generate two perturbed copies of the target data, randomly publish one of them, and then compare the model’s membership score on the published vs. the unpublished one. While their framework can work for different domains such as foundation models, they still require the data holders to mark a substantial portion (1$\sim$10%) of dataset. Instead, MembershipTracker considers the more challenging (and realistic) setting where the marked data amounts to $\leq 0.1\%$ of the dataset. We also “stress test” their technique under a similar setting as ours and find that their performance degrades considerably then (see Appendix F).
3 Problem Formulation
@@
Membership Inference (MI) Game. We denote ${F}_{\theta}:\mathcal{X}\rightarrow[0,1]^{n}$ as a model that maps an input sample $x\in\mathcal{X}$ to a probability vector over $n$ classes. $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ is a training set sampled from some distribution $\mathcal{D}$. ${F}_{\theta}\leftarrow\mathcal{T}(D)$ denotes a model ${F}_{\theta}$ produced from running the training algorithm $\mathcal{T}$ on $D$. Next, we define a MI game, which proceeds between a challenger $\mathcal{C}$ and an adversary $\mathcal{A}$.
  1. 1.
@@
    The challenger samples a dataset $D\leftarrow\mathcal{D}$, and a target point $z\leftarrow\pi$ (such that $D\cap\pi=\emptyset$).
  2. 2.
@@
  4. 4.
    The adversary emits a guess $\hat{z}\in\pi$.
  5. 5.
@@
For a specific target point $x$, $\pi=\{x,\perp\}$, where $\perp$ indicates the absence of an example. The adversary’s goal is to guess whether the model $F$ is trained on $D$ or $D\cup\{x\}$.
3.1 MI for Data Provenance
@@
  1. 1.
    The user is allowed to modify ${X}\in\pi$ ($\pi=\{{X},\perp\}$).
  2. 2.
    The challenger samples a dataset $D\leftarrow\mathcal{D}$, and a target set $Z\leftarrow\pi$ (such that $D\cap\pi=\emptyset$).
  3. 3.
@@
  5. 5.
    The user emits a guess $\hat{Z}\in\pi$.
  6. 6.
@@
For a given user $U$, $\pi=\{X,\perp\}$, where $\perp$ indicates the absence of the target samples by $U$, and the goal is to determine whether $F$ is trained on $D$ or $D\cup\{X\}$.
3.1.1 User capability

### v2

@@ Abstract.
This work proposes MembershipTracker, a practical data auditing tool that can empower ordinary users to reliably detect the unauthorized use of their data in training DL models. We view data auditing through the lens of membership inference (MI). MembershipTracker consists of a lightweight data marking component to mark the target data with small and targeted changes, which can be strongly memorized by the model trained on them; and a specialized MI-based verification process to audit whether the model exhibits strong memorization on the target samples.
@@
MembershipTracker only requires the users to mark a small fraction of data (0.005$\sim$0.1in proportion to the training set), and it enables the users to reliably detect the unauthorized use of their data (average 0% FPR@100% TPR). We show that MembershipTracker is highly effective across various settings, including industry-scale training on the full-size ImageNet-1k dataset. We finally evaluate MembershipTracker under multiple classes of countermeasures.
1. Introduction
@@
Modern deep learning (DL) models have attained remarkable performance in various tasks such as image classification and language generation. Their success is largely driven by the availability of massive training data (Rombach et al., 2022; Ramesh et al., 2022). However, it is not always clear whether the data used to train these models has been used with the permission of the data owners. For instance, Clearview.ai, in developing their facial recognition system, scraped billions of user photos from social media platform without the users’ consent (Hill, 2020). The unregulated use of personal data for building DL models has engendered mounting concerns (Dixit, 2023; Vincent, 2023; Hill, 2020), and is in violation of privacy laws such as the European Union’s General Data Protection Regulation (GDPR) (gdp, [n. d.]). In spite of this, data holders often have limited agency in detecting the unauthorized use of their data due to the lack of practical data provenance tools.
@@
This leads to two main lines of work for data provenance. The first line of work seek to inject designated features into the user data, which can induce the model to exhibit a certain detectable behavior for provenance purposes (Sablayrolles et al., 2020; Hu et al., 2022; Li et al., 2022a; Wenger et al., 2024). The injected features can take the form of a backdoor trigger pattern (Hu et al., 2022; Li et al., 2022a); or a spurious feature pattern that can cause output probability shift by the model (Wenger et al., 2024). However, these techniques either impose strong assumptions on the users (e.g., the ability to mark a large fraction of data) (Sablayrolles et al., 2020; Li et al., 2022a; Hu et al., 2022) and/or suffer from limited auditing performance (Sablayrolles et al., 2020; Li et al., 2022a; Wenger et al., 2024).
@@
The second line of work is based on membership inference (MI). There are instance- and user-level MI - the former infers whether a specific sample is used to train a model (Shokri et al., 2017; Carlini et al., 2022); while the latter detects the usage of any of the user’s data (Song and Shmatikov, 2019; Chen et al., 2023b; Kandpal et al., 2023). Our work focuses on auditing whether specific user samples are used for training the target model. Unfortunately, existing MI methods often suffer from limited effectiveness, as measured by their true positive rate (TPR) and false positive rate (FPR) (Carlini et al., 2022). However, a high-power MI method is needed to facilitate data provenance with low FPR (to avoid false accusation) and high TPR (for correct data provenance).
@@
To improve, recent studies propose several techniques based on data poisoning (Tramèr et al., 2022; Chen et al., 2022; Chaudhari et al., 2023). Their idea is to inject mislabeled samples into the model’s training set, which can amplify the model’s behavioral changes on some target samples, thereby making them easier to be de-identified. To apply these techniques for data auditing, users have to be able to mislabel some training data, which, unfortunately, excludes the real-world scenarios where the data labels are assigned by the model owner or an external service (Dzieza, 2023; Deng et al., 2009). This renders these techniques challenging for real-world use.
@@
Figure 1. MembershipTracker is a data provenance tool that operates by: (1) marking the target data (e.g., facial images, artworks) with small and targeted changes, and then (2) initiating a specialized membership inference process to audit whether the target data are used for training the model.
@@
MembershipTracker operates under the realistic setting where we assume the users have neither the capability nor expertise to train shadow/proxy models (Ye et al., 2022; Bertran et al., 2024; Carlini et al., 2022), or to manipulate the data labels (Tramèr et al., 2022), as these are well beyond the capacity of ordinary personal users (Section 3.1.1). Rather, MembershipTracker only requires the users to mark a small fraction of data (amounting to 0.005$\sim$0.1of the training set) for provenance purposes. Our goal is to empower the users to carry out high-power MI on their target data with high TPR and low FPR, given black-box access to the model.
Technical design. We first observe that the effectiveness of MI on a given model is directly related to the model’s ability in memorizing individual data points, and hence samples that are strongly memorized by the model (e.g., outlier samples) are easier to be de-identified (Yeom et al., 2018; Carlini et al., 2022; Tramèr et al., 2022; Chen and Pattabiraman, 2024a). Therefore, in order to fulfill our goal, there are two challenges as follows.
  1. (1)
@@
For the first challenge, we propose a lightweight data marking technique to mark the target data with small and targeted changes, which are formulated as the combination of outlier feature and procedural noise (Co et al., 2019). These subtle changes can be readily created by the users without assuming access to a proxy model, and can still induce the model to strongly memorize the specially-marked samples while preserving their visual information (e.g., Fig. 1).
For the second challenge, given a target user’s small set of marked samples, we propose a novel set-based MI process to audit whether they are used to train the model. Our approach follows a common idea in existing user-level MI studies (Song and Shmatikov, 2019; Meeus et al., 2024; Kandpal et al., 2023) to leverage the collective information across the small set of target samples for MI. But unlike prior methods, MembershipTracker does not require any additional shadow/reference models (Song and Shmatikov, 2019; Meeus et al., 2024; Kandpal et al., 2023) and can enable high-power MI for reliable data provenance.
@@
We find that by merely marking a small fraction of samples (0.005$\sim$0.1of the dataset), MembershipTracker effectively empowers the users to reliably audit the usage of their data (average 0% FPR@100% TPR), and it is scalable to the large-scale ImageNet training (on both Convolutional and Transformer network). To the best of our knowledge, MembershipTracker is the first technique that can scale to support reliable data auditing in large-scale training setting. This renders MembershipTracker a practical data auditing tool that contributes to responsible AI practice.
2. Related Work
Our work focuses on detecting the unauthorized use of personal data in training DL models, which is complementary to existing work that aim to protect the data from unwanted use by rendering them un-learnable (Shan et al., 2020; Huang et al., 2021). Hence we focus on existing literature for data provenance, and we classify them into two categories.
@@ 2. Related Work
Dataset level provenance aims to detect whether a DL model is trained on a specific dataset (Maini et al., 2020; Sablayrolles et al., 2020; Li et al., 2022a). Maini et al. propose dataset inference, which detects whether two models trained on the same dataset exhibit similarities in their decision boundaries (Maini et al., 2020). Other work modify portions of training set to leave a certain artifact on the models, such as radioactive data (Sablayrolles et al., 2020) and backdoor watermark (Li et al., 2022a). However, recent work (Huang et al., 2024) finds that these techniques (Sablayrolles et al., 2020; Li et al., 2022a) still have poor auditing performance. Moreover, dataset level solutions require several assumptions that are outside the scope of our work. First, many of them require control over a nontrivial portion of training set (e.g., 10$\sim$20(Sablayrolles et al., 2020; Li et al., 2022a)). Some solutions assume access to some known feature extractor (Sablayrolles et al., 2020), access to the training set or the ability to train a proxy model (Maini et al., 2020; Sablayrolles et al., 2020). These are well beyond the user capability we consider (in Section 3.1.1).
User level provenance seeks to detect whether the users’ data are used to train a DL model (Hu et al., 2022; Wenger et al., 2024), and they require only modifying the users’ own data. Their idea is to mark the users’ data with a designated feature to induce a detectable behavior into the model. Hu et al. propose to inject backdoor trigger feature into the user data (Hu et al., 2022), but they assume the user capability to manipulate the data labels. Other studies on clean-label backdoor attacks may be repurposed to overcome this, but they still require training a proxy model to compute the poisoned data (Souri et al., 2022; Zeng et al., 2023).
Wenger et al. propose to inject a target spurious feature into the user data to be learned by the suspected model, and then detect whether the model has associated the injected feature with the target class label (Wenger et al., 2024). To detect, the users can separately overlay the target and non-target features to some auxiliary data (outside the target class), and then determine whether the target feature causes a slightly higher prediction probability on the target class, compared with the non-target features (e.g., 10% vs. 1%). Both their work and ours involve adding spurious features to mark users’ data, but there are several major differences between the two. Notably, our analysis reveals that they (Wenger et al., 2024) significantly underestimate their detection false positives (from 0% to >30%). This is because they employ a discrepant detection procedure for the testing features that are used for training (true positives), and those that are not (false positives); however, a fair evaluation should follow a consistent procedure (see our detailed analysis in Appendix D).
@@
Instance-level MI detects whether a specific instance is used to train a model. It was first proposed by Shokri et al. (Shokri et al., 2017) and there are many follow-up studies (Yeom et al., 2018; Choquette-Choo et al., 2021; Bertran et al., 2024; Ye et al., 2022; Carlini et al., 2022). However, even state-of-the-art MI methods (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022) still struggle with achieving high TPR when controlled at the low FPR regime (more in Section 4.1).
@@
This has led to several attempts at improving the effectiveness of existing MI methods via data poisoning. Tramer et al. propose to inject mislabeled samples into the model’s training set, which can transform the target samples into outliers and amplify their influence on the model’s decision, thereby making the target samples easier to be de-identified (Tramèr et al., 2022). While injecting mislabeled samples represents an effective solution for improving the MI success (Tramèr et al., 2022; Chen et al., 2022; Chaudhari et al., 2023), it may be challenging for the users to apply to their data, because in many scenarios the users may not have control over the data labels (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Even if this is feasible, we find that such a method can still only increase the MI success to a limited extent (validated in Appendix E). By contrast, MembershipTracker does not assume control of the data labels, and is still able to facilitate high-power MI for tracing data provenance.
@@
User-level MI aims to detect whether any of a user’s data is used to train a model. Unlike instance-level MI, the auditor possesses a set of samples from a target user that are not necessarily used to train the target model. Song et al. (Song and Shmatikov, 2019) propose the first user-level MI for natural language models, which first trains multiple shadow models, and then uses the outputs from shadow models as the features to train an audit model. Subsequent work develop solutions for other settings, including speech recognition models (Miao et al., 2021; Chen et al., 2023a), metric learning models (Li et al., 2022b; Chen et al., 2023b) and large language models (Kandpal et al., 2023; Meeus et al., 2024).
@@
A common thread of existing user-level MI methods is to aggregate the information across multiple samples (e.g., different voices of the same speaker) to perform MI (Song and Shmatikov, 2019; Chen et al., 2023a, b; Kandpal et al., 2023). The MI process in MembershipTracker follows a similar philosophy. However, prior work often requires access to additional reference models (Kandpal et al., 2023) or shadow models (Song and Shmatikov, 2019; Chen et al., 2023b, a), and they still have limited MI effectiveness (e.g., limited TPR under the low FPR regime (Kandpal et al., 2023; Meeus et al., 2024)). MembershipTracker overcomes these limitations by means of a novel data-marking and specialized MI auditing process, which can simultaneously achieve high TPR and low FPR for data provenance.
@@
Recent work by Huang et al. (Huang et al., 2024) proposes a general framework for auditing unauthorized data use in training DL models. Their idea is to generate two perturbed copies of the target data, randomly publish one of them, and then compare the model’s membership score on the published vs. the unpublished one. While their framework can work for different domains such as foundation models, they still require the data holders to mark a substantial portion (1$\sim$10%) of dataset. Instead, MembershipTracker considers the more challenging (and realistic) setting where the marked data amounts to $\leq 0.1\%$ of the dataset. We also “stress test” their technique under a similar setting as ours and find that their performance degrades considerably then (see Appendix F).
3. Problem Formulation
@@
Membership Inference (MI) Game. We denote ${F}_{\theta}:\mathcal{X}\rightarrow[0,1]^{n}$ as a model that maps an input sample $x\in\mathcal{X}$ to a probability vector over $n$ classes. $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ is a training set sampled from some distribution $\mathcal{D}$. ${F}_{\theta}\leftarrow\mathcal{T}(D)$ denotes a model ${F}_{\theta}$ produced from running the training algorithm $\mathcal{T}$ on $D$.
Next, we define a MI game, which proceeds between a challenger $\mathcal{C}$ and an adversary $\mathcal{A}$. $\mathcal{A}$ has to guess which element from some universe $\mathcal{U}$ was used to train a model. For a specific target point $x$, $\mathcal{U}=\{x,\perp\}$, where $\perp$ indicates the absence of an example.
  1. (1)
@@
    The challenger samples a dataset $D\leftarrow\mathcal{D}$, and a target point $z\leftarrow\mathcal{U}$ (such that $D\cap\mathcal{U}=\emptyset$).
  2. (2)
@@
  4. (4)
    The adversary emits a guess $\hat{z}\in\mathcal{U}$.
  5. (5)
@@
3.1. MI for Data Provenance
@@
  1. (1)
    The user is allowed to modify ${X}\in\mathcal{U}$ ($\mathcal{U}=\{{X},\perp\}$).
  2. (2)
    The challenger samples a dataset $D\leftarrow\mathcal{D}$, and a target set $Z\leftarrow\mathcal{U}$ (such that $D\cap\mathcal{U}=\emptyset$).
  3. (3)
@@
  5. (5)
    The user emits a guess $\hat{Z}\in\mathcal{U}$.
  6. (6)
@@
For a given user $U$, $\mathcal{U}=\{X,\perp\}$, where $\perp$ indicates the absence of the target samples by $U$, and the goal is to determine whether $F$ is trained on $D$ or $D\cup\{X\}$.
3.1.1. User capability